The rate constants of a reaction at 500K and 700K are 0.02s–1 and 0.07s–1 respectively. Calculate the values of Ea and A.

Grade:8

1 Answers

Sunil Kumar FP
9 years ago
It is given by the formula
log(k2/k1)=Ea/2.303R(1/T1 -1/T2)
log(.07/.02)=Ea/2.303*8.314(1/500-1/700)
Ea=18228kJ
now again we have
k=Ae^(-Ea/RT)
.02=Ae^(-18228/8.314*500)
A=1.6s^-1
The above are the values of Ea and A